OpenStudy (anonymous):

find the unit vector in the direction of vector {7,-24}

OpenStudy (anonymous):

The unit vector will have fractions for the x and y components of the unit vector. The numerators will be the same as for the given vector. The denominator for each of the x and y will be the length of the given vector.

OpenStudy (anonymous):

So, can you get the length of the given vector? Think of the vecotr as starting from the origin, a basic vector. What is the length from (0, 0) to (7, -24)? Use the distance formula for 2 points. That's the denominator.

OpenStudy (anonymous):

When you do this, you will get: (7/d)^2 + (-24/d)^2 = 1^2 = 1

OpenStudy (anonymous):

@comput313 ? Is this making sense now?

OpenStudy (anonymous):

yes i think so, let me try to work it out.

OpenStudy (anonymous):

so the unit vector will be {7/31,-24/31}?

OpenStudy (anonymous):

Not sure where you got the 31, possibly a typo somewhere in your calculaions. Try: (7)^2 + (-24)^2 = d^2

OpenStudy (anonymous):

That's from: [0 - 7]^2 + [0 - (-24)]^2 = d^2

OpenStudy (anonymous):

shouldnt that be -7^2 then? 0-7=-7?

OpenStudy (anonymous):

that dosent matter does it.

OpenStudy (anonymous):

(-7)^2 will be the same as (7)^2 because you can have either: (0 - 7)^2 or (7 - 0)^2 The order will not matter, and the distance will always be positive because it is a measurable physical concept, not implying direction like vectors. The direction of the given vector and the unit vector will be suggested by the coordinates, basically by the "numerators".

OpenStudy (anonymous):

so it comes out to d=25

OpenStudy (anonymous):

Yes! Good job! @comput313

OpenStudy (anonymous):

so the unit vector is {7/25,-24/25}?

OpenStudy (anonymous):

There you go! Once you have the unit vecotr concept down, the direction is the same as the given vecotr. It is just shortened or lengthened.

OpenStudy (anonymous):

how could this be written in terms of i and j?

OpenStudy (anonymous):

For a vector {a, b} you put it into: ai + bj The "i" and "j" are usually in bold face in books.

OpenStudy (anonymous):

ok, thanks!

OpenStudy (anonymous):

Good luck to you in all of your studies and thx for the recognition! @comput313 And you're welcome!